OpenStudy (anonymous):

when 2 less than a number n is increased by70% the result is8.5

OpenStudy (tkhunny):

We have the number, n. Do what it says. 2 less than n How do you write that. Increase the result by 70% How do you write that?

OpenStudy (acxbox22):

1.7(n-2)=8.5 n-2=5 n=7

OpenStudy (tkhunny):

@acxbox22 Please read the Code of Conduct and heed its wisdom.

OpenStudy (acxbox22):

sorry, should have left out the n=7 part lol

OpenStudy (tkhunny):

Should have provided explanations and allowed the OP to show and do some work.

OpenStudy (acxbox22):

then i would have "guided the user to the answer

OpenStudy (acxbox22):

*"

OpenStudy (anonymous):

\[(n-2)+\left( n-2 \right)\times \frac{ 70 }{ 100 }=8.5\] calculate n